# ChatTS: Aligning Time Series with LLMs via Synthetic Data for Enhanced Understanding and Reasoning

### v2

@@ 4.1.5. Implementation
For GPT-based models, we used OpenAI’s API to infer and track token consumption. For ChatTS and QWen-based models, the training and inference are conducted locally on 8$\times$×A800 GPUs. The token consumption for ChatTS is calculated after the “Reorder & Concat” step. Full-parameter training is used for ChatTS with DeepSpeed [^6] https://www.deepspeed.ai/ and LLaMA-Factory (Zheng et al., 2024), with Qwen2.5-14B-Instruct (Yang et al., 2024) [^7] https://huggingface.co/Qwen/Qwen2.5-14B-Instruct as the base model. Inference for both Qwen and ChatTS is also conducted with DeepSpeed.
@@
The evaluation results on alignment tasks are shown in Table 3. It can be observed that ChatTS achieves leading performance in almost all tasks and datasets. Overall, ChatTS achieves 46.0%–82.6% improvement in categorical metrics and 80.7%–153.1% in numerical metrics. This indicates that compared to current industry-leading models such as GPT-4o, our ChatTS achieves significant superiority in overall time series alignment capability. Furthermore, this also demonstrates that even if only synthetic data are used, it is possible to achieve good alignment results with real-world time series.
At the same time, we find that the GPT-4o (Vision) model achieves the best result among the baselines. This suggests that existing vision-modal MLLMs have a basic capability to analyze the overall shape characteristics of time series. However, regarding numerical capabilities and detailed analysis (e.g., noise analysis), vision-based models perform slightly worse than text-based models. This is likely because LLMs have a certain degree of numerical computation ability, enabling text-based methods to perform calculations based on the numbers in text format. In contrast, vision-based methods may have limited accuracy in recognizing detailed information from images due to limited image resolution. Refer to Section 5 for a more detailed analysis. Compared with Text/Vision/Agent-based methods, ChatTS features a native multimodal time series capability, directly accepting time series arrays as input. This enables both global and detailed time series information to be preserved. Therefore, through large-scale alignment training, our model surpasses all the baseline models in global and local feature analyses.
@@
In multivariate tasks, ChatTS demonstrates more advantages. For text-based models, the encoded input prompt becomes excessively long, requiring a very large context length, making it difficult to summarize the multivariate information accurately. For vision-based models, multivariate time series are plotted simultaneously on a single image, making it difficult to identify features from different time series accurately. In contrast, ChatTS is equipped with a context-aware time series encoding, which helps accurately analyze the referenced time series based on contextual information.
Agent-based models performed poorly across all tasks compared with vision-based methods. This may be attributed to insufficient tool accuracy and overly long CoT, preventing it from accurately integrating the time series attributes. Refer to Section 4.6 for a more detailed analysis of agent-based models.
Additionally, we compared the number of input tokens consumed by different models and the estimated input token costs (the inference cost of ChatTS is estimated based on the official API inference pricing for QWen2.5). The results are presented in Table 3. It can be observed that, due to ChatTS’s use of a native multimodal time series encoding, it requires only a minimal number of tokens to encode time series data, resulting in costs significantly lower than those of the baseline models (as low as 13.4% of the GPT-4o-mini model) while outperforming the baselines in terms of effectiveness. This indicates that ChatTS has sufficient competitiveness in real-world time series analysis applications.
@@
4.4. RQ3. Study of Time Series Modality
To investigate the effectiveness of the time series multimodality in ChatTS, we performed an ablation study based on a text-only version of ChatTS (w/o TS Modality). We remove the time series encoder in ChatTS (i.e. using the original QWen-2.5 model) and use the same training data with ChatTS (the time series arrays are encoded into text) in model training. The experimental results are shown in Figure 9 and Figure 10. Overall, the model using only the text modality performs significantly worse than the original ChatTS model. This indicates that encoding multimodal information is crucial for accurately capturing both shape and numerical information. However, in certain sub-evaluation metrics (e.g., noise), the text-only model outperforms the multimodal ChatTS, suggesting that text modality models still have strong capabilities for identifying small fluctuations. In MTS tasks, the text-only model is nearly incapable of answering any questions. This implies that even with extensive multivariate training data, text-only LLMs still struggle to handle multivariate problems due to excessively long context lengths because of severe hallucinations and inaccurate responses. This further demonstrates the effectiveness of using time series MLLMs in multivariate tasks.
4.5. RQ4. Ablation Studies on Synthetic Training Data
@@
To further explore the effectiveness of attribute-based time series + text data generation and TSEvol, we conducted the following ablation studies: 1) w/o Attribute-Based. All of the training datasets (PT + SFT) are replaced by the GPT-generated datasets provided in  (Merrill et al., 2024). This dataset was directly generated using GPT-produced Python code. Corresponding Q&As are also generated by GPT according to the code. By replacing all the training data (generated with the attribute-based generator), we can identify the contribution of the attribute-based data generation method to the training results. 2) w/o TSEvol. The SFT datasets (generated with TSEvol) are replaced with an SFT dataset directly generated using an LLM. TSEvol was not used to generate this dataset, but prompts were crafted to encourage the LLM to generate diverse QA pairs. For fairness, we also incorporated the instruct-following dataset into the training set for these ablation studies to ensure the IF capability of the models.
The evaluation results of the retrained models are shown in Figure 9 and Figure 10. It can be observed that the model trained on GPT-generated data performed significantly worse across various alignment evaluation tasks compared to ChatTS, in particular for tasks related to local fluctuation and numerical analysis. This is likely due to discrepancies between the GPT-generated time series data and the corresponding captions, which fail to describe local feature details and specific values accurately. Furthermore, we observed that models trained with TSEvol achieved significant improvements in reasoning capabilities compared to other models, along with more minor improvements in alignment tasks. This indicates that TSEvol effectively diversifies the forms of questions and formulates Q&As targeting different time series attributes, enabling the model to achieve better alignment and reasoning performance.
4.6. RQ5. Study of Agent-Based Methods
Figure 11. Performance of the Agent (GPT-4o-mini) Model Based on “Perfect Tools”. We implement tools with different accuracies based on feature labels and show the performance of the Agent-based model under these tools with differing accuracies. The Agent-based model exhibits a high sensitivity to the accuracy of the tools.
Agent-based methods are widely used in current applications. However, in the evaluations for RQ1 and RQ2, we found that the performance of Agent-Based models fell short of expectations. Two primary issues were identified: (1) the tools invoked by Agent models lacked accuracy, and (2) the output format of Agent models did not meet the required specifications, leading to parsing failures.
To further investigate the performance upper bound of Agent-Based models, we conducted a more detailed analysis:
@@
    Using attribute labels from the synthetic evaluation data, we developed a set of “perfect tools” with strictly controlled accuracy (by randomly using incorrect labels). Therefore, these tools rely on time series labels and are not applicable in real-world scenarios but are restricted to experimental investigations.
@@
    We removed responses from the Agent model that failed to parse, ensuring all responses were valid.
Eliminating the influence of tool implementation and formatting issues allows us to analyze the performance potential of Agent-Based models precisely. We conducted a sensitivity analysis of the Agent-Based model (GPT-4o-mini) using the “perfect tools” with varying accuracy levels. The experimental results shown in Figure 11 reveal that the Agent-Based model is highly sensitive to tool accuracy within the range of [0.9, 1.0]. When the tools achieve perfect accuracy (accuracy = 1.0), the Agent-Based model slightly outperforms ChatTS in numerical metrics. However, its scores in numerical metrics and inductive reasoning tasks remain significantly lower than ChatTS. A critical factor contributing to this performance gap is their poor performance in multivariate tasks. In such cases, an increased number of tool calls and longer CoT tend to occur, leading to hallucinations and inaccuracies.
Thus, even with high-accuracy tools, the performance of Agent-Based models remains extremely sensitive to minor tool errors. Furthermore, in the analysis of multivariate time series, which involves multi-step reasoning and summarization, the overall capability of Agent-Based models is heavily constrained by the reasoning and summarization limitations of the underlying LLM. This significantly restricts their performance. In contrast, ChatTS has native multimodal time series capabilities, enabling it to analyze multiple time series simultaneously, effectively reducing reasoning complexity and improving accuracy.
@@ 5.1. Real-World Application: DB Operation
Figure 12. An application case of ChatTS in a failure diagnosis with an Oracle database system. The “¡ts¿¡ts/¿” tags indicate the insertion points of time series arrays.
Although the application of time series + LLMs has been fully demonstrated in some work (Zhou et al., 2023a; Wang et al., 2024), their implementations typically heavily rely on agent tools. To illustrate the performance of ChatTS with its native time series multimodal capability in real-world applications, we present a typical Oracle DB operation application through a multivariate time series-related multi-turn dialogue with ChatTS.
In this case study, we show a scenario where an Oracle DB operator has identified a recent anomaly and retrieves several time series metrics from the monitoring system, inputting them into ChatTS for analysis (as shown in Figure 12) [^8] The input of time series data can be achieved in various ways, such as by using APIs to feed the collected metrics into ChatTS. The system implementation details are out of the scope of this paper.. By querying ChatTS, the operator obtains the names of all metrics with anomalies near the points of interest. Subsequently, to accurately pinpoint the root cause, the operator provides ChatTS with a textual document titled “Oracle Database Troubleshooting Rulebook” and requests ChatTS to analyze the root cause and propagation of the system failure step-by-step, combining insights from the rulebook and the time series anomalies. Notably, the rulebook is entirely in text form, without a strictly structured format, which is helpful for the operators to share their expert experience effectively.
In this case, ChatTS can accurately identify anomalies and amplitudes in multivariate time series. By leveraging “the metric with the largest fluctuation” in the rulebook, ChatTS can further reason about the root cause and failure propagation path. This further shows that ChatTS can effectively utilize its alignment capability to analyze time series and perform complex analysis in real-world applications with its robust reasoning ability.
5.2. Detailed Time Series Analysis
@@ 5.3. Baseline Comparison
(a) Case Study on Noise and Local Fluctuation
@@
To further demonstrate the advantages of the native time series modality in ChatTS, we present several case studies on the comparison with vision/text-based methods in alignment tasks.
5.3.1. Noise
Identifying noise requires precise global understanding and numerical reasoning about time series. The noise level is in conjunction with the global patterns to determine whether it is significant relative to overall variations. In Figure 14(a), we show the results of a question about noise analysis with different models. It can be observed that although these models can identify the presence of noise in the time series, the vision- and text-based models incorrectly classify the noise level, considering the noise is minor overall compared to the global variance. In contrast, ChatTS shows superior performance over the baseline models, accurately identifying both the noise level and its std.
5.3.2. Local Fluctuations
Analysis of local fluctuations requires models to have accurate capability to capture subtle changes of shape or numerical features in time series. In Figure 14(a), both ChatTS and the text-based models accurately identified the sudden fluctuation and its amplitude. However, while correctly identifying this change, the vision-based models exhibited significant amplitude deviation. Thus, it can be observed that the vision-based models struggled to accurately extract the specific numerical values before and after the sudden increase from the visual representation, leading to an incorrect estimation of the amplitude. This highlights that, compared to using images as inputs, directly leveraging the native time series as input can better preserve the numerical information.
5.3.3. Trend
Trend analysis primarily relies on the ability to analyze the overall shape of time series data. In Figure 14(b), we present a misleading case for trend analysis. In this example, the original image can easily lead one to conclude that the overall trend of the time series is steady. However, when the two sharp upward spikes are ignored (see the cropped plot), we can easily find a significant decreasing trend in this time series.
As expected, the vision-based model makes a mistake in this case, incorrectly concluding that the time series remained steady. Furthermore, its numerical responses on the amplitude of change are highly inaccurate. This demonstrates that vision-based models are sensitive to image scaling, impairing their ability to accurately analyze overall shape and numerical features. Similarly, while correctly identifying the starting point’s value, the text-based model lacked the capability to understand the global shape of the time series. In contrast, ChatTS, with its ability of native time series encoding, accurately captured both the overall shape and numerical features of the time series.
@@ 6. Related Work
Time Series Question Answering (TSQA). With the rapid development of LLMs, TSQA systems have combined the reasoning capabilities of LLMs with time series analysis to enable more efficient cross-domain decision-making and complex task handling (Jin et al., 2024). Time series question-answering systems have been explored in various fields, such as AIOps (Zhou et al., 2023a; Wang et al., 2024), IoT (Xing et al., 2021; Gallo et al., 2023), healthcare (Yu et al., 2023; Oh et al., 2024), finance (Maître et al., 2020; Kurisinkel et al., 2024), and traffic (Da et al., 2024; Lai et al., 2023). However, these methods are often limited to agent-based (Yao et al., 2022) and retrieval-augmented generation (RAG) (Lewis et al., 2020) approaches, lacking a comprehensive understanding of time series and sufficient reasoning capabilities. Although some recent studies (Chow et al., 2024) have attempted to leverage temporal multimodal approaches for time series reasoning tasks, they typically rely on task-specific corpora. They are trained and evaluated on specific tasks (e.g., classification tasks or forecasting tasks), lacking multivariate analysis capabilities. Compared to the research on multimodal question answering in fields like images and videos, time series question answering still lacks robust multimodal alignment methods and evaluation frameworks (Merrill et al., 2024). Therefore, in contrast to existing studies, this paper is the first to propose a comprehensive time series modality alignment and fine-tuning process, evaluated using multiple alignment and reasoning tasks.
LLM + Time Series. In addition to the research above, many studies have combined LLMs with time series for various downstream tasks, leveraging the powerful capabilities of LLMs (Jin et al., 2023; Gruver et al., 2024; Su et al., 2024; Chang et al., 2023; Liu et al., 2024b; Cao et al., 2023; Zhou et al., 2023b). However, while these models are using LLMs as backbones, they are designed for specific downstream tasks and lack language alignment capabilities, making them unsuitable for question answering and reasoning applications. Moreover, some studies employ vision-based multimodal LLMs for time series prediction (Chen et al., 2024) and anomaly detection (Zhuang et al., 2024). This approach aligns with the vision-based LLM methods discussed in this paper but is significantly constrained in its ability to analyze time series.

### v3

@@ 4.1.5. Implementation
For GPT-based models, we used OpenAI’s API to infer and track token consumption. For ChatTS and QWen-based models, the training and inference are conducted locally on 8$\times$×A800 GPUs. The token consumption for ChatTS is calculated after the “Reorder & Concat” step.
@@
The evaluation results on alignment tasks are shown in Table 3. ChatTS consistently outperforms all baseline models across nearly all tasks and datasets, achieving 46.0%–75.9% improvement in categorical metrics and 80.7%–112.7% in numerical metrics compared to industry-leading models like GPT-4o. This demonstrates that synthetic training data can effectively enable strong alignment with real-world time series.
Among the baselines, GPT-4o (Vision) performs best, suggesting vision-based MLLMs possess some capability to analyze shape characteristics of time series, though they remain limited by image resolution when interpreting details. Text-based methods struggle with the constraints of prompt length, while agent-based approaches performed below expectations (see Section 4.7 for detailed analysis).
@@
ChatTS’s advantages are particularly pronounced in multivariate tasks, where text-based models face challenges with excessively long prompts and vision-based models struggle to distinguish features across multiple time series plotted simultaneously. In contrast, ChatTS’s context-aware time series encoding accurately analyzes referenced time series based on contextual information.
From the efficiency perspective, ChatTS’s native multimodal encoding requires significantly fewer tokens to represent time series data, resulting in much lower costs compared with the baselines (see Table 3). This shows both the effectiveness and efficiency of treating time series as a native modality.
@@
4.4. RQ3. Studies of Synthetic Training Data
@@
To evaluate the effectiveness of attribute-based time series generation and TSEvol, we conducted ablation studies with two variants: (1) w/o Attribute-Based, where all training datasets were replaced by GPT-generated datasets from (Merrill et al., 2024), containing time series directly generated using GPT-produced Python code with corresponding GPT-generated Q&As; (2) w/o TSEvol, where SFT datasets were replaced with data directly generated using an LLM without the evolutionary approach, though with prompts designed to encourage diversity. Both variants included the instruct-following dataset to ensure fair comparison.
The evaluation results in Figures 9 and 10 reveal that models trained on GPT-generated data performed significantly worse across alignment tasks, particularly for local fluctuation detection and numerical analysis. This suggests the attribute-based generation method better captures precise feature details and numerical values. Meanwhile, models trained with TSEvol demonstrated substantial improvements in reasoning capabilities and modest gains in alignment tasks, indicating that TSEvol effectively diversifies question formats and generates tailored Q&As for different time series attributes, enhancing overall model performance.
4.5. RQ4. Scaling of Training Dataset
(a) Train Data
(b) SFT on Text-Based LLMs
Figure 11. Scaling of training dataset and text-based LLMs.
Figure 11(a) illustrates the relationship between ChatTS performance and training data size. The results show that increasing the Phase 1 training dataset size from 10% to 100% of the current size significantly improves performance, but further expansion yields minimal gains. Thus, our chosen training set size is well-balanced, ensuring sufficient data for effective alignment while avoiding too much resource consumption during training.
4.6. RQ5. Study of Time Series Modality
To investigate the effectiveness of the time series multimodality in ChatTS, we performed an ablation study based on a text-only version of ChatTS (w/o TS Modality). We remove the time series encoder in ChatTS (i.e. using the original QWen-2.5 model) and use the same training data with ChatTS (the time series arrays are encoded into text) in model training. The experimental results are shown in Figure 9 and Figure 10. Overall, the model using only the text modality performs significantly worse than the original ChatTS model. This indicates that encoding multimodal information is crucial for accurately capturing both shape and numerical information. However, in certain sub-evaluation metrics (e.g., noise), the text-only model outperforms the multimodal ChatTS, suggesting that text modality models still have strong capabilities for identifying small fluctuations. In MTS tasks, the text-only model is nearly incapable of answering any questions. This implies that even with extensive multivariate training data, text-only LLMs still struggle to handle multivariate problems due to excessively long context lengths because of severe hallucinations and inaccurate responses. Additionally, to compare the performance gains between text-based LLMs and TS-MLLMs, we fine-tuned various sizes of the Qwen2.5 series text-based LLMs using the text version of the ChatTS training dataset (as shown in Figure 11(b)). Experimental results indicate that even when fine-tuning the larger Qwen2.5-32B text model, the results still do not outperform those of ChatTS (14B), which has native multimodal capabilities. This further validates the importance of native multimodal capabilities in ChatTS, whether in the accuracy for MTS analysis or cost efficiency (see Table 3).
4.7. RQ6. Study of Agent-Based Methods
(a) Agent with different tool accuracy.
(b) Agent w/o different tools.
(c) Study of error cases.
Figure 12. Agent with “Perfect Tools”. Even with perfect tools, Agent still makes errors (e.g., Error Tool Using)
Agent-based methods are widely applied but showed suboptimal performance in our evaluations (RQ1, RQ2) due to several main issues: (1) tool inaccuracy (2) error tool use, and (3) response formatting that caused parsing failures. To explore their performance upper bound, we conducted detailed analyses:
@@
    Parsing Failures: We exclude responses that failed to parse, ensuring all outputs were valid.
@@
    Perfect Tools: We design “perfect tools” with controlled accuracy through time series labels in the synthetic dataset (Figure 12(a)). The accuracy of the perfect tools can be strictly controlled.
  3. (3)
    Tool Ablation: We perform ablation studies (Figure 12(b)) to evaluate the impact of individual tools on accuracy.
  4. (4)
    Error Analysis: We categorize agent errors into three types: Error Tool Using, Misunderstanding, and Hallucination. We analyze their impact on performance (Figure 12(c)).
Our sensitivity analysis (Figure 12(a)) shows that agent performance is highly sensitive to tool accuracy, especially in the [0.9, 1.0] range. For categorical and numerical tasks, the agent with perfect tools slightly outperforms ChatTS on UTS tasks but lags behind in MTS tasks. For agents, MTS tasks typically require more tool calls and reasoning, which places higher demands on LLMs’ tool-using and summarization capabilities. In contrast, ChatTS processes multiple time series natively, reducing complexity and improving accuracy for MTS tasks. The tool ablation study (Figure 12(b)) shows that Agent depends heavily on both tool precision and completeness, particularly for MTS tasks. Even with all tools available, agents frequently fail to invoke the correct tool at the right time (e.g., using the classification tool rather than the anomaly detection tool to identify the position of a spike), limiting their effectiveness. Error analysis (Figure 12(c)) reveals “Error Tool Using” as the largest source of errors. When these cases are excluded, agent accuracy exceeds 95%, surpassing ChatTS. This validates the correctness of the implementation of perfect tools and the model, which also shows their limitations: Agents may struggle with tool selection and reasoning.
In summary, while perfect tools improve agent performance, challenges such as tool selection errors, misunderstandings, and hallucinations persist, leaving agents less effective than ChatTS for complex time series tasks.
@@ 5. Case Studies and Applications
5.1. Case Studies on Real-World Data
Figure 13. Case studies on real-world time series data.
To investigate the performance of ChatTS on real-world time series, we perform several case studies with challenging questions, the results are shown in Figure 13.
5.1.1. Shape and Statistical Analysis
The “Basic Shape Analysis” case demonstrates ChatTS’s capability to analyze an NYC taxi passenger time series with complex periodic fluctuations and local anomalies. ChatTS accurately identifies multiple trend segments, the periodicity, and the upward spike along with their amplitudes. This shows ChatTS’s capability to capture both global patterns and localized features. In the “Statistic Analysis” case, ChatTS analyzes advertisement CPC data with misleading scaling. Despite potential confusion in the minimum value, ChatTS correctly identifies the max/min values and their positions. These cases show ChatTS’s robustness in statistical analysis in complex real-world time series.
5.1.2. OOD Fluctuation Recognition
The “OOD Fluctuation” case presents ChatTS with a traffic occupancy time series containing an OOD fluctuation pattern absent from its training data. However, ChatTS accurately describes it as a “Convex-Shaped Elevation”, characterized by a gradual rise followed by a sharper decline along with the overall shape of a convex. This demonstrates ChatTS’s inherent understanding of time series patterns themselves, rather than simply repeating representations from the training set. This indicates that ChatTS has a certain capacity to generalize to real-world data despite being trained exclusively on synthetic data.
@@ 5.2. Real-World Application: DB Operation
Figure 14. An application case of ChatTS in a failure diagnosis with an Oracle database system.
To illustrate the performance of ChatTS with its native time series multimodal capability in real-world applications, we present a typical Oracle DB operation application through an MTS-related multi-turn dialogue with ChatTS. In this case study, an Oracle DB operator has identified a recent anomaly and retrieves several time series metrics from the monitoring system, inputting them into ChatTS for analysis (as shown in Figure 14) [^2] DB Metrics can be input into ChatTS through API. The system implementation details are out of the scope of this paper.. By querying ChatTS, the operator obtains the names of all metrics with anomalies. Then, to accurately pinpoint the root cause, the operator provides ChatTS with a textual document titled “Oracle Database Troubleshooting Rulebook” and requests ChatTS to analyze the root cause and propagation of the system failure step-by-step, combining insights from the rulebook and the time series anomalies. Notably, the rulebook is entirely in text form, without a strictly structured format, which is helpful for the operators to share their expert experience effectively. The responses of ChatTS show that it can accurately identify anomalies and amplitudes in multivariate time series. By leveraging “the metric with the largest fluctuation” in the rulebook, ChatTS can further reason about the root cause and failure propagation path. This further shows that ChatTS can effectively utilize its alignment capability to analyze time series and perform complex analysis in real-world applications with its robust reasoning ability.
5.3. Real-World Application: Detailed Analysis
@@ 5.4. Baseline Comparison
(a) Case Study on Complex Time Series Caption
@@
5.4.1. Seasonality, Trend and Fluctuations
As shown in Figure 16(a), due to tool inaccuracy, the Agent fails to identify the periodicity. This further led the Agent to misinterpret periodic patterns as different trend changes, resulting in errors in trend analysis. Moreover, the LLM does not realize the problem or attempt to correct it. Similarly, the Vision-based model also exhibited errors in analyzing local fluctuations and periodicity. In contrast, ChatTS, with its time-series modality awareness, accurately captures the periodicity and trend transitions. This case shows a key limitation of agent-based tools: the precision of tools alone cannot overcome the cascading errors caused by the initial error of time series patterns.
5.4.2. Detailed Trend Analysis
Figure 16(b) presents a misleading case where the original image suggests a steady trend. However, in the cropped plot that ignores the two spikes, there is a significant decreasing trend. Due to the subtle nature of trends displayed in the time series images, the Vision-based model incorrectly classified them as steady. Similarly, text-based models, while identifying starting values, fail to identify the global shape of the time series. In contrast, ChatTS captures both the overall trend and numerical details accurately due to its native time series encoding capabilities.
@@ 6. Related Work
Time Series Question Answering (TSQA). With the rapid development of LLMs, TSQA systems have combined the reasoning capabilities of LLMs with time series analysis to enable more efficient cross-domain decision-making and complex task handling (Jin et al., 2024). Time series question-answering systems have been explored in various fields, such as AIOps (Zhou et al., 2023a; Wang et al., 2024), IoT (Xing et al., 2021; Gallo et al., 2023), healthcare (Yu et al., 2023; Oh et al., 2024), finance (Maître et al., 2020; Kurisinkel et al., 2024), and traffic (Da et al., 2024; Lai et al., 2023). However, these methods are often limited to agent-based (Yao et al., 2022) and retrieval-augmented generation (RAG) (Lewis et al., 2020) approaches, lacking a comprehensive understanding of time series and sufficient reasoning capabilities. Although some recent studies (Chow et al., 2024) have attempted to leverage temporal multimodal approaches for time series reasoning tasks, they typically rely on task-specific corpora. They are trained and evaluated on specific tasks (e.g., classification tasks or forecasting tasks), lacking multivariate analysis capabilities. Compared to the research on multimodal question answering in fields like images and videos, time series question answering still lacks robust multimodal alignment methods and evaluation frameworks (Merrill et al., 2024; Cai et al., 2024). Therefore, in contrast to existing studies, this paper is the first to propose a comprehensive time series modality alignment and fine-tuning process, evaluated using multiple alignment and reasoning tasks.
LLM + Time Series. In addition to the research above, many studies have combined LLMs with time series for various downstream tasks, leveraging the powerful capabilities of LLMs (Jin et al., 2023; Gruver et al., 2024; Su et al., 2024; Chang et al., 2023; Liu et al., 2024b; Cao et al., 2023; Zhou et al., 2023b; Cai et al., 2023). However, while these models are using LLMs as backbones, they are designed for specific downstream tasks and lack language alignment capabilities, making them unsuitable for question answering and reasoning applications. Moreover, some studies employ vision-based multimodal LLMs for time series prediction (Chen et al., 2024) and anomaly detection (Zhuang et al., 2024). This approach aligns with the vision-based LLM methods discussed in this paper but is significantly constrained in its ability to analyze time series.